the second no. of three nos.is 3 more than twice the first no. The 3rd no.is two times the 2nd no.If the difference between 3 times the 2nd no. and five times the 1st no. is 18.What is the 3rd no.?
Answer by ankor@dixie-net.com(22740)

3 numbers, a, b, c
:
Write an equation for each statement:
:
"the second no. of three nos.is 3 more than twice the first no."
b = 2a + 3
:
"The 3rd no.is two times the 2nd no."
c = 2b
:
"If the difference between 3 times the 2nd no. and five times the 1st no. is 18."
3b - 5a = 18
or
5a = 3b - 18
Replace b with (2a+3) from the 1st equation
5a = 3(2a+3) - 18
5a = 6a + 9 - 18
5a = 6a - 9
9 = 6a - 5a
a = 9
Find b using the 1st equation
b = 2(9) + 3
b = 18 - 3
b = 21
find c using the 2nd equation
c = 2(21)
c = 42, is the 3rd number